# Communication Cost Reduction for Subgraph Counting under Local Differential Privacy via Hash Functions

### v1

@@ V-D Communication Cost
The upload cost of user $v_{i}$ in our mechanism is $\mathcal{O}\left(\log{\frac{n}{s}}\left(\frac{d_{i}}{s}+\frac{n}{s^{2}(e^{\varepsilon_{1}}-1)}\right)\right)$.
@@
Recall that, in the randomized response step, each user $v_{i}$ submit the vector $[a^{\prime}_{i,c_{1}},\dots,a^{\prime}_{i,c_{m}}]\in\{0,1\}^{m}$ for $m\approx n/s$. To save the communication cost, the user can send the index set $\{c_{t}:a^{\prime}_{i,c_{t}}=1\}\subseteq\{1,\dots,m\}$. If $\ell$ elements of the vector are one, the cost is $\ell\cdot\log m\approx\ell\cdot\log n/s$.
Now, we discuss the expected value of $\ell$. As each $a^{\prime}_{i,j}=1$ is sampled in the vector with probability $1/s$ and $e^{\varepsilon^{\prime}}-1=s(e^{\varepsilon_{1}}-1)$, we have that
$\mathbb{E}[\ell]$ $=$ $\left(\frac{e^{\varepsilon^{\prime}}}{1+e^{\varepsilon^{\prime}}}d_{i}+\frac{1}{1+e^{\varepsilon^{\prime}}}(n-d_{i})\right)/s$
$\leq$ $\frac{d_{i}}{s}+\frac{1}{s^{2}(e^{\varepsilon_{1}}-1)}(n-d_{i})$
$=$ $O\left(\frac{d_{i}}{s}+\frac{n}{s^{2}(e^{\varepsilon_{1}}-1)}\right).$
@@
The download cost for all users is $\mathcal{O}\left(\log{\frac{n}{s}}\left(\frac{|E|}{s}+\frac{n}{s^{2}(e^{\varepsilon_{1}}-1)}\right)\right)$.
@@
As users have to download information from all others, the download cost is then:
$\sum_{i}\mathcal{O}\left(\log{\frac{n}{s}}\left(\frac{d_{i}}{s}+\frac{n}{s^{2}(e^{\varepsilon_{1}}-1)}\right)\right)$
$=$ $\mathcal{O}\left(\log{\frac{n}{s}}\left(\frac{|E|}{s}+\frac{n^{2}}{s^{2}(e^{\varepsilon_{1}}-1)}\right)\right).$
VI Application: Triangle Counting
While our GRR mechanism is adaptable for counting various subgraphs, the prevalent focus in existing literature, including [20, 22, 24], has been on triangle counting under local differential privacy. Therefore, in this section, we focus on triangle counting as well, enabling a direct comparison with these studies.
VI-A Sensitivity and Loss of Our Mechanism
Recall from the previous section that, in the counting step of our mechanism, we ask each user $v_{i}$ to estimate $|\Delta_{i}|$ when $\Delta_{i}=\{(v_{i},v_{j},v_{k}):i<j<k,\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{j},v_{k}\}\in E\}$. The number of triangles in the graph $G$, denoted by $|\Delta_{G}|$ is then the sum of all $|\Delta_{i}|$. The estimation of $|\Delta_{i}|$ in our mechanism is $\tilde{\Delta}_{i}:=\sum\limits_{j,k:\{i,j\},\{i,k\}\in E}\tilde{a}_{j,h_{j}(k)}$ added with the Laplacian noise when the variable $\tilde{a}_{j,h_{j}(k)}$ is defined in the previous section.
The magnitude of the Laplacian noise, dictated by the sensitivity of $\tilde{\Delta}_{i}$, is capped at $d_{i}\cdot\omega_{i}$, where $d_{i}$ is the degree of node $v_{i}$ and $\omega_{i}$ is the unbiased factor outlined in the previous section. If the degree $d_{i}$ is publicly known, we can directly set the sensitivity to $d_{i}\cdot\omega_{i}$. In scenarios where $d_{i}$ is not public, the sensitivity must be adjusted to the maximum degree, $d_{\text{max}}=\max_{i}d_{i}$. This maximum degree can be estimated as $\tilde{d}_{\text{max}}=C+\max_{i}\tilde{d}_{i}$, with $C$ being a positive integer and $\tilde{d}_{i}$ representing the degree announced in the degree sharing step. It is important to note, however, that this publication could potentially underestimate the actual maximum degree, leading to some bias in our publication. If a node $i$ has a degree $d_{i}$ that exceeds $\tilde{d}_{\text{max}}$, it is required to eliminate edges from its local data.
From now, let us suppose that $d_{i}$ is publicly known and $Lap_{i}$ is a noise sampled from the Laplacian distribution with parameter $\omega_{i}\cdot d_{i}/\varepsilon_{2}$. Consequently, the estimation of user $v_{i}$ is $\tilde{\Delta}_{i}+Lap_{i}$. In the preceding section, we established that $\tilde{\Delta}_{i}+Lap_{i}$ serves as an unbiased estimate of $|\Delta_{i}|$. The primary source of estimation loss arises from the variance of $\sum_{i}(\tilde{\Delta}_{i}+Lap_{i})$, which we consider in the subsequent theorem.
@@ VI-A Sensitivity and Loss of Our Mechanism
Let $S_{2}$, $C_{4}$, and $W_{4}$ are the number of two-star, four-cycle, and walks of length four in the graph $G$. The variance of our estimation $\sum_{i}(\tilde{\Delta}_{i}+Lap_{i})$ is $\mathcal{O}\left(\left(\frac{s}{(e^{\varepsilon_{1}}-1)\varepsilon_{2}}\right)^{2}\cdot S_{2}+\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}\times(S_{2}+C_{4}+\frac{s}{n}\cdot W_{4})\right)$.
@@
The variance of $\sum\limits_{i}\left(\sum\limits_{j,k:\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E}\tilde{a}_{j,h_{j}(k)}+Lap_{i}\right)$ comes from four parts:
@@
    $\sum_{i}\text{Var}\left(Lap_{i}\right)$: By the parameter of the Laplacian noise, we know that the variance is $O\left(\sum_{i}\left(\frac{s\cdot d_{i}}{(e^{\varepsilon_{1}}-1)\varepsilon_{2}}\right)^{2}\right)=O\left(\left(\frac{s}{(e^{\varepsilon_{1}}-1)\varepsilon_{2}}\right)^{2}\cdot S_{2}\right)$;
@@
    $\sum\limits_{i}\left(\sum\limits_{j,k:\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E}\text{Var}\left(\tilde{a}_{j,h_{j}(k)}\right)\right)$: We have from Theorem 4 that $\text{Var}\left(\tilde{a}_{j,h_{j}(k)}\right)$ is $O\left(\frac{s}{(e^{e_{1}}-1)^{2}}\right)$. As there are $S_{2}$ terms in the summation, we have that the summation is no more than $O\left(\frac{s\cdot S_{2}}{(e^{e_{1}}-1)^{2}}\right)$.
@@
    The covariance resulting from the repeated counting of $\tilde{a}_{j,h_{j}(k)}$ in the summation: We observe that $\tilde{a}_{j,h_{j}(k)}$ might be counted twice if there exist $i,i^{\prime}>j>k$ such that $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{i^{\prime}},v_{j}\},\{i^{\prime},k\}\in E$. From Theorem 4, the covariance for each instance of double counting is equal to $\text{Var}\left(\tilde{a}_{j,h_{j}(k)}\right)=O\left(\frac{s}{(e^{e_{1}}-1)^{2}}\right)$, and the maximum frequency of such double counting is limited to no more than $C_{4}$.
  4. 4.
    The covariance between the terms $\tilde{a}_{j,h_{j}(k)}$ and $\tilde{a}_{j,h_{j}(k^{\prime})}$ arises in situations where both terms are included in the summation: This occurs when there are $i>j,k$ and $i^{\prime}>j,k^{\prime}$ such that the edge sets $\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{i^{\prime}},v_{j}\},\{v_{i^{\prime}},v_{k^{\prime}}\}$ are part of $E$. These instances occur no more than $W_{4}$ times. According to Theorem 5, each occurrence contributes to the covariance with a value of $\frac{s^{2}}{n}\cdot\frac{1}{(e^{e_{1}}-1)^{2}}$.
VI-B Further Optimization for Triangle Counting: Central Server Sampling
@@
VI-B1 Modification
Remember that through the GRR mechanism, the server acquires the set $\bar{\Delta}_{i}=\{t\mid a^{\prime}_{i,c_{t}}=1\}$ from user $v_{i}$ following the randomized response step. This set is then shared with all users at the start of the counting step. As identified in the prior subsection, this approach not only leads to significant communication costs but also contributes to a substantial covariance in triangle counting. To address these issues, we suggest generating subsets $\bar{\Delta}_{i}^{(1)},\dots,\bar{\Delta}_{i}^{(n)}$ from $\bar{\Delta}_{i}$ independently, and then distributing the subset $\bar{\Delta}_{i}^{(j)}$ to user $v_{j}$. An element $t$ in $\bar{\Delta}_{i}$ is independently selected for inclusion in $\bar{\Delta}_{i}^{(j)}$ with a constant probability, which we denote as $\mu_{c}$.
Let $a^{(i)}_{j,h_{j}(k)}\in\{0,1\}$ is a random variable indicating if $h_{j}(k)\in\bar{\Delta}_{j}^{(i)}$. The user $v_{i}$ believe that there is an edge $\{v_{j},v_{k}\}$ in the graph $G$ if $a^{(i)}_{j,h_{j}(k)}=1$. We have that $\mathbb{E}\left[a^{(i)}_{j,h_{j}(k)}\right]=\mu_{c}\cdot\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}\right]$ for all $i,j,k$.
Recall the variables $\omega_{i}$ and $\tilde{\sigma}_{i}$ defined in the previous section. We define $\tilde{a}^{(i)}_{j,h_{j}(k)}=\frac{\omega_{i}}{\mu_{c}}a^{(i)}_{j,h_{j}(k)}+\tilde{\sigma}_{i}$.
@@ VI-B1 Modification
By the definitions and Theorem 3, we have that
$\mathbb{E}[\tilde{a}^{(i)}_{j,h_{j}(k)}]$ $=$ $\frac{\omega_{i}}{\mu_{c}}\mathbb{E}[a^{(i)}_{j,h_{j}(k)}]+\mathbb{E}[\tilde{\sigma}_{i}]$
$=$ $\omega_{i}\mathbb{E}[a^{\prime}_{j,h_{j}(k)}]+\mathbb{E}[\tilde{\sigma}_{i}]$
$=$ $a_{j,k}.$
From the previous theorem, our unbiased estimation of the number of triangles under this modification is then $\sum\limits_{i}\left(\sum\limits_{j,k:\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E}\tilde{a}^{(i)}_{j,h_{j}(k)}+Lap_{i}\right)$ when $Lap_{i}$ is a noise sampled from the Laplacian distribution with parameter $d_{i}\cdot\omega_{i}/(\mu_{c}\cdot\varepsilon_{2})$.
VI-B2 Variance of the Modified Mechanism
We first consider the variance of $\tilde{a}^{(i)}_{j,h_{j}(k)}$ in the subsequent theorem:
@@ VI-B2 Variance of the Modified Mechanism
[…]
@@
Because $a^{(i)}_{j,h_{j}(k)}$ is a binomial variable, we have from Lemma 1 that:
$\text{Var}\left(a^{(i)}_{j,h_{j}(k)}\right)$
$=$ $\mathbb{P}\left(a^{(i)}_{j,h_{j}(k)}=1\right)\cdot\left(1-\mathbb{P}\left(a^{(i)}_{j,h_{j}(k)}=1\right)\right)$
$=$ $\mu_{c}\cdot\mathbb{P}\left({a^{\prime}}_{j,h_{j}(k)}=1\right)\cdot\left(1-\mu_{c}\cdot\mathbb{P}\left(a^{(i)}_{j,h_{j}(k)}=1\right)\right)$
$\leq$ $\mu_{c}\cdot\mathbb{P}\left({a^{\prime}}_{j,h_{j}(k)}=1\right)$
$\leq$ $\mu_{c}\frac{1+\sigma_{j}}{\omega_{j}}.$
Then, we have from Lemma 2 that:
$\text{Var}\left(\tilde{a}^{(i)}_{j,h_{j}(k)}\right)$ $=$ $\frac{\omega_{j}^{2}}{\mu_{c}^{2}}\text{Var}\left(a^{(i)}_{j,h_{j}(k)}\right)+\text{Var}\left(\tilde{\sigma}_{i}\right)$
$\leq$ $\frac{1}{\mu_{c}}\frac{1+\sigma_{j}}{\omega_{j}}+\frac{s^{2}}{n^{2}\varepsilon_{0}^{2}}$
$=$ $O\left(\frac{s}{\mu_{c}(e^{\varepsilon_{1}}-1)^{2}}+\frac{s^{2}}{n^{2}\varepsilon_{0}^{2}}\right).$
At first glance, when contrasting this outcome with the findings in Theorem 4, it might appear that employing central server sampling amplifies the variance of the estimations, potentially diminishing their utility. However, it is important to recognize that for a communication cost reduction by $1/\mu_{c}$ times. Additionally, in the same way as the four-cycle trick in Section III, the central server sampling can reduce the covariance stated in Theorem 7. In particular, it significantly reduces the covariance for the graph with a lot of walks with length four (or graph with large number of $W_{4}$).
@@
The covariance of the estimators $\tilde{a}^{(i)}_{j,h_{j}(k)}$ and $\tilde{a}^{(i^{\prime})}_{j^{\prime},h_{j}^{\prime}(k^{\prime})}$ is $O\left(\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}\right)$ when $j=j^{\prime}$, $i\neq i^{\prime}$ and $k=k^{\prime}$, $O\left(\frac{s^{2}}{n(e^{\varepsilon_{1}}-1)^{2}}\right)$ when $j=j^{\prime}$, $i\neq i^{\prime}$ and $k,k^{\prime}<j$, and $O\left(\frac{s^{2}}{\mu_{c}n(e^{\varepsilon_{1}}-1)^{2}}\right)$ when $j=j^{\prime}$, $i=i^{\prime}$ and $k,k^{\prime}<j$. It is equal to zero in other cases.

### v2

@@ 5.5 Communication Cost
The upload cost of user $v_{i}$ in our mechanism is $O\left(\left(\frac{d_{i}}{s}+\frac{n}{s^{2}(e^{\varepsilon_{1}}-1)}\right)\log{\frac{n}{s}}\right).$
@@
Recall that, in the randomized response step, each user $v_{i}$ submits a binary vector $a^{\prime}_{i,c_{1}},\dots,a^{\prime}_{i,c_{m}}]\in\{0,1\}^{m}$, where $m\approx n/s$. To reduce communication cost, the user may instead transmit only the index set $\{t:a^{\prime}_{i,c_{t}}=1\}\subseteq\{1,\dots,m\}$. If $\ell$ entries in the vector are equal to one, the communication cost is $\ell\cdot\log m\approx\ell\cdot\log(n/s)$.
We now analyze the expected value of $\ell$. Since each $a_{i,j}$ is sampled into the vector with probability $1/s$, and noting that $e^{\varepsilon^{\prime}}-1=s(e^{\varepsilon_{1}}-1)$, we obtain:
$\mathbb{E}[\ell]=\left(\frac{e^{\varepsilon^{\prime}}}{1+e^{\varepsilon^{\prime}}}\cdot d_{i}+\frac{1}{1+e^{\varepsilon^{\prime}}}\cdot(n-d_{i})\right)\cdot\frac{1}{s}\leq\frac{d_{i}}{s}+\frac{1}{s^{2}(e^{\varepsilon_{1}}-1)}\cdot(n-d_{i})=O\left(\frac{d_{i}}{s}+\frac{n}{s^{2}(e^{\varepsilon_{1}}-1)}\right).$
@@
The download cost for all users is $O\left(\log{\frac{n}{s}}\left(\frac{|E|}{s}+\frac{n^{2}}{s^{2}(e^{\varepsilon_{1}}-1)}\right)\right).$
@@
Since each user must download information from all other users, the total download cost is
$\sum_{i}O\left(\log\left(\frac{n}{s}\right)\left(\frac{d_{i}}{s}+\frac{n}{s^{2}(e^{\varepsilon_{1}}-1)}\right)\right)=O\left(\log\left(\frac{n}{s}\right)\left(\frac{|E|}{s}+\frac{n^{2}}{s^{2}(e^{\varepsilon_{1}}-1)}\right)\right),$
where $|E|=\sum_{i}d_{i}$ is the total number of edges. ∎
When $|E|=O(n)$, the upload cost is $O\left(\frac{n}{s^{2}}\log n\right)$ and the download cost is $O\left(\frac{n^{2}}{s^{2}}\log n\right)$. Recall that the communication cost in the two-step mechanism is $\Theta(n^{2})$. Therefore, our technique reduces the communication cost by a factor of $s^{2}$.
6 Use Case: Triangle Counting
While our GroupRR mechanism is applicable to the counting of various subgraphs, the majority of prior work—such as Imola et al. (2021; 2022a); Eden et al. (2023)—has focused specifically on triangle counting under local differential privacy. To facilitate direct comparison with these existing approaches, we also focus on triangle counting in this section.
In Subsection 6.1, we analyze the estimation loss introduced by our mechanism, excluding the effect of Line 5 in the counting step of Algorithm 3, where noise is added via either the smooth sensitivity or local Laplace mechanism. In Subsection 6.2, we propose an algorithmic refinement aimed at further reducing this estimation loss for triangle counting. Finally, Subsections 6.3 and 6.4 are devoted to analyzing the noise contribution introduced in the final step of the algorithm.
6.1 Sensitivity and Loss of Our Mechanism
In the counting step of our mechanism, each user $v_{i}$ estimates $|\Delta_{i}|$ when $\Delta_{i}=\{(v_{i},v_{j},v_{k}):k<j<i,\{v_{i},v_{j}\},\{v_{i},v_{k}\},\{v_{j},v_{k}\}\in E\}$. The number of triangles in the graph $G$, denoted by $|\Delta_{G}|$ is then the sum of all $|\Delta_{i}|$. The estimation of $|\Delta_{i}|$ in our mechanism is $|\tilde{\Delta}_{i}|:=\sum\limits_{j,k:\{i,j\},\{i,k\}\in E}\tilde{a}_{j,h_{j}(k)}$ added with the Laplacian noise when the variable $\tilde{a}_{j,h_{j}(k)}$ is defined in the previous section. The variance of our mechanism without the noise in the final step of our mechanism is shown in the subsequent theorem.
@@ 6.1 Sensitivity and Loss of Our Mechanism
Let $S_{2}$, $C_{4}$, and $W_{4}$ denote the number of two-stars, four-cycles, and walks of length four in the graph $G$, respectively. Then, the variance of the estimator $\sum_{i}\sum\limits_{{j,k:\\ \{i,j\},\{i,k\}\in E}}\tilde{a}_{j,h_{j}(k)}$ is bounded by
$O\left(\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}\cdot\left(S_{2}+C_{4}+\frac{s}{n}\cdot W_{4}\right)\right).$
@@
The variance of the estimator $\sum_{i}\sum\limits_{{j,k:\\ \{i,j\},\{i,k\}\in E}}\tilde{a}_{j,h_{j}(k)}$ can be decomposed into the following three components:
@@
    Variance of individual terms: From Theorem 4, we have $\text{Var}\left(\tilde{a}_{j,h_{j}(k)}\right)=O\left(\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}\right)$. As there are $S_{2}$ such terms (one for each two-star), the total contribution from the variances is
    $O\left(\frac{s\cdot S_{2}}{(e^{\varepsilon_{1}}-1)^{2}}\right).$
@@
    Covariance due to repeated terms: A term $\tilde{a}_{j,h_{j}(k)}$ may appear multiple times in the summation if there exist $i,i^{\prime}>j>k$ such that $\{i,j\},\{i,k\},\{i^{\prime},j\},\{i^{\prime},k\}\in E$. Such configurations correspond to four-cycles, of which there are at most $C_{4}$. From Theorem 4, each such repeated term contributes $O\left(\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}\right)$ to the total covariance, yielding a contribution bounded by
    $O\left(\frac{s\cdot C_{4}}{(e^{\varepsilon_{1}}-1)^{2}}\right).$
@@
    Covariance between distinct terms: Dependencies may also arise between distinct terms $\tilde{a}_{j,h_{j}(k)}$ and $\tilde{a}_{j,h_{j}(k^{\prime})}$ if there exist $i>j,k$ and $i^{\prime}>j,k^{\prime}$ such that the edge sets $\{i,j\},\{i,k\},\{i^{\prime},j\},\{i^{\prime},k^{\prime}\}\in E$. These configurations correspond to walks of length four involving shared center nodes. There are at most $W_{4}$ such configurations, and each contributes $O\left(\frac{s^{2}}{n(e^{\varepsilon_{1}}-1)^{2}}\right)$ to the total covariance, resulting in an overall contribution of
    $O\left(\frac{s^{2}\cdot W_{4}}{n(e^{\varepsilon_{1}}-1)^{2}}\right).$
Combining all three components, we conclude that the total variance is
$O\left(\frac{s}{(e^{\varepsilon_{1}}-1)^{2}}\cdot\left(S_{2}+C_{4}+\frac{s}{n}\cdot W_{4}\right)\right).$
6.2 Further Optimization for Triangle Counting: Central Server Sampling
@@
6.2.1 Modification
Remember that through the GroupRR mechanism, the server acquires the set $\bar{\xi}_{i}=\{t\mid a^{\prime}_{i,c_{t}}=1\}$ from user $v_{i}$ following the group randomized response step. This set is then shared with all users at the start of the counting step. As identified in the prior subsection, this approach not only leads to significant communication costs but also contributes to a substantial covariance in triangle counting. To address these issues, we suggest generating subsets $\bar{\xi}_{i}^{(1)},\dots,\bar{\xi}_{i}^{(n)}$ from $\bar{\xi}_{i}$ independently, and then distributing the subset $\bar{\xi}_{i}^{(j)}$ to user $v_{j}$. An element $t$ in $\bar{\xi}_{i}$ is independently selected for inclusion in $\bar{\xi}_{i}^{(j)}$ with a constant probability, which we denote as $\mu_{c}$. The central server sampling is described in Algorithm 4.
1
2Function CentralSampling
 Input: For all user $j$, the grouped and obfuscated adjacency list $(a^{\prime}_{j,c_{1}},\dots,a^{\prime}_{j,c_{m}})$ obtained from the randomized response step, central server sampling probability $\mu_{C}$
 Output: For all user $i,j$, the grouped and obfuscated adjacency list for user $j$, which user $i$ will use for counting the number of triangles in the counting step, $(a^{(i)}_{j,c_{1}}\dots,a^{(i)}_{j,c_{m}})$.
3 [Server] For each $i,j,t$, $a_{j,c_{t}}^{(i)}=0$ if $a^{\prime}_{j,c_{t}}=0$. If $a^{\prime}_{j,c_{t}}=1$, $a_{j,c_{t}}^{(i)}=1$ with probability $\mu_{c}$, and $a_{j,c_{t}}^{(i)}=0$ with probability $1-\mu_{c}$.;
4 [Server] Send $(a^{(i)}_{j,c_{1}}\dots,a^{(i)}_{j,c_{m}})$ to user $i$.
5
Algorithm 4 Central Server Sampling: This process is executed between Step 3 and 4 of the GroupRR mechanism
Let $a^{(i)}_{j,h_{j}(k)}\in\{0,1\}$ is a random variable indicating if $h_{j}(k)\in\bar{\xi}_{j}^{(i)}$. The user $v_{i}$ believe that there is an edge $\{v_{j},v_{k}\}$ in the graph $G$ if $a^{(i)}_{j,h_{j}(k)}=1$. We have that $\mathbb{E}\left[a^{(i)}_{j,h_{j}(k)}\right]=\mu_{c}\cdot\mathbb{E}\left[a^{\prime}_{j,h_{j}(k)}\right]$ for all $i,j,k$.
Recall the variables $\omega_{i}$ and $\tilde{\sigma}_{i}$ defined in the previous section. We define $\tilde{a}^{(i)}_{j,h_{j}(k)}=\frac{\omega_{i}}{\mu_{c}}a^{(i)}_{j,h_{j}(k)}-\tilde{\sigma}_{i}$.
@@ 6.2.1 Modification
By the definitions and Theorem 3, we have
$\mathbb{E}[\tilde{a}^{(i)}_{j,h_{j}(k)}]=\frac{\omega_{i}}{\mu_{c}}\cdot\mathbb{E}[a^{(i)}_{j,h_{j}(k)}]-\mathbb{E}[\tilde{\sigma}_{i}]=\omega_{i}\cdot\mathbb{E}[a^{\prime}_{j,h_{j}(k)}]-\mathbb{E}[\tilde{\sigma}_{i}]=a_{j,k}.$
From the previous theorem, our unbiased estimation of the number of triangles under this modification is then $|\tilde{\Delta}_{i}|=\sum\limits_{j,k:\{v_{i},v_{j}\},\{v_{i},v_{k}\}\in E}\tilde{a}^{(i)}_{j,h_{j}(k)}$ added by the Laplacian noise or the noise from the smooth sensitivity mechanism.
6.2.2 Variance of the Modified Mechanism
We first consider the variance of $\tilde{a}^{(i)}_{j,h_{j}(k)}$ in the subsequent theorem.
@@ 6.2.2 Variance of the Modified Mechanism
$\text{Var}\left(\tilde{a}^{(i)}_{j,h_{j}(k)}\right)=O\left(\frac{s}{\mu_{c}(e^{\varepsilon_{1}}-1)^{2}}+\frac{s^{2}}{n^{2}\varepsilon_{0}^{2}}\right)$.
@@
Since $a^{(i)}_{j,h_{j}(k)}$ is a binomial random variable, Lemma 4 implies
$\text{Var}\left(a^{(i)}_{j,h_{j}(k)}\right)$ $=$ $\Pr\left[a^{(i)}_{j,h_{j}(k)}=1\right]\cdot\left(1-\Pr\left[a^{(i)}_{j,h_{j}(k)}=1\right]\right)$
$=$ $\mu_{c}\cdot\Pr\left[a^{\prime}_{j,h_{j}(k)}=1\right]\cdot\left(1-\mu_{c}\cdot\Pr\left[a^{\prime}_{j,h_{j}(k)}=1\right]\right)$
$\leq$ $\mu_{c}\cdot\Pr\left[a^{\prime}_{j,h_{j}(k)}=1\right]\leq\mu_{c}\cdot\frac{1+\sigma_{j}}{\omega_{j}}.$
Then, by Lemma 5, we have
$\text{Var}\left(\tilde{a}^{(i)}_{j,h_{j}(k)}\right)=\frac{\omega_{j}^{2}}{\mu_{c}^{2}}\cdot\text{Var}\left(a^{(i)}_{j,h_{j}(k)}\right)+\text{Var}\left(\tilde{\sigma}_{i}\right)\leq\frac{1}{\mu_{c}}\cdot\frac{1+\sigma_{j}}{\omega_{j}}+\frac{s^{2}}{n^{2}\varepsilon_{0}^{2}}.$
Therefore,
$\text{Var}\left(\tilde{a}^{(i)}_{j,h_{j}(k)}\right)=O\left(\frac{s}{\mu_{c}(e^{\varepsilon_{1}}-1)^{2}}+\frac{s^{2}}{n^{2}\varepsilon_{0}^{2}}\right).$
At first glance, when contrasting this outcome with the findings in Theorem 4, it might appear that employing central server sampling amplifies the variance of the estimations, potentially diminishing their utility. However, it is important to recognize that for a communication cost reduction by $1/\mu_{c}$ times. Additionally, in the same way as the four-cycle trick in Section 3, the central server sampling can reduce the covariance stated in Theorem 7. In particular, it significantly reduces the covariance for the graph with a lot of walks with length four (or graph with large number of $W_{4}$).
@@
The covariance of $\tilde{a}^{(i)}_{j,h_{j}(k)}$ and $\tilde{a}^{(i^{\prime})}_{j^{\prime},h_{j}^{\prime}(k^{\prime})}$ is $O\left(s/(e^{\varepsilon_{1}}-1)^{2}\right)$ when $j=j^{\prime}$, $i\neq i^{\prime}$ and $k=k^{\prime}$, $O\left(s^{2}/(n(e^{\varepsilon_{1}}-1)^{2})\right)$ when $j=j^{\prime}$, $i\neq i^{\prime}$ and $k,k^{\prime}<j$, and $O\left(s^{2}/(\mu_{c}n(e^{\varepsilon_{1}}-1)^{2})\right)$ when $j=j^{\prime}$, $i=i^{\prime}$ and $k,k^{\prime}<j$. It is equal to zero in other cases.